OpenStudy (anonymous):

Find the equation of the tangent through the external point (8,-1) to the ellipse 2x^2+5y^2=70

OpenStudy (anonymous):

help..

OpenStudy (anonymous):

@ganeshie8

OpenStudy (loser66):

Take derivative of the curve, then plug it into the tangent line equation.

OpenStudy (loser66):

4x + 10yy'=0 hence \(y'= \dfrac{-4x}{10y}\), at (8,-1), \(y'= 16/5\)

OpenStudy (anonymous):

4x+10y(dy/dx)=0 4(8)+10(-1)dy/dx=0?

OpenStudy (anonymous):

ok then wats next?

OpenStudy (loser66):

that is m, right? that is the slope of the tangent line at (8,-1), y - y_0 = m (x -x_0) , that is it.

OpenStudy (loser66):

where \((x_0, y_0) =(8,-1)\)

ganeshie8 (ganeshie8):

we need to be careful, (8, -1) is not on the ellipse

OpenStudy (loser66):

Wow!! ok, let me check

OpenStudy (loser66):

oh, yea, so, the tangent line to the curve which passes through (8,-1) right? @ganeshie8

OpenStudy (anonymous):

ok so y+1=16/5x-128/5 y=(16/5)x+128/5

OpenStudy (loser66):

Sorry AliLnn. I misread the question. hehehe... I guided you wrong until the derivative. Let @ganeshie8 take over. @ganeshie8 Please.

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):